Calculate the % sodium in a 0.721 g detergent sample. 

Final yield is 0.320 g NaP2O4


1
Expert's answer
2021-02-09T03:44:34-0500

The mass of sodium in 0.320 g of NaP2O4 is the number of the moles of NaP2O4 times the molar mass of sodium:

"n(NaP_2O_4) = n(Na)"

"n(NaP_2O_4) = \frac{m(NaP_2O_4) }{M(NaP_2O_4) } = \frac{0.721 }{148.93 } = 0.00484" mol.

"m(Na) = n(Na) \u00b7M(Na) = 0.00484\u00b722.99"

"m(Na) = 0.111" g.

The mass percentage of the sodium in the sample is:

"w = \frac{m(Na)}{m(sample)} ·100\%= \frac{0.111}{0.721} ·100\% = 15.4\%" .

Answer: the % sodium in the detergent sample is 15.4%.